OpenStudy (anonymous):

Anyone help me with volumes matter please....? How to find the volume of a solid S of a right circular cone with height h and base radius r?

OpenStudy (anonymous):

What I try is.. as it was found that \[r = \sqrt{c ^{2}-h ^{2}}\] so the cross sectional area is A(h)= pi r^2 = pi (c^2-h^2) from the drawing.. a=0, b=h so the volume is \[\int\limits_{0}^{h}A(h) dh = \int\limits_{0}^{h} pi*(c ^{2}-h ^{2}) dh = 1/3 pi * r ^{2} h\] but this wasn't quite right.

OpenStudy (anonymous):

@candyfloss94 u r in wrong place u must in somewhere else.. i send u link go there and put ur question there . this is computer science room ...

OpenStudy (anonymous):

yeah, I know, sorry about that. That's why I close this question.

OpenStudy (anonymous):

http://openstudy.com/study#/groups/Mathematics

OpenStudy (anonymous):

@candyfloss94 don't be sorry i am also in mathematics rum but i don't try any question but there are so many in there . :) :)